Given Ca=20 c=MPC=3/4 Ip=I=20 a; determine equilibrium level of income when there is no government sector b;define equilibrium level of income When government spending is equal to 25 &no taxation c; define equilibrium level of income When government expenditure and tax is equal to 25

1
Expert's answer
2022-02-04T06:36:33-0500

a) if no government

Y= C+I

C= a+ bY

= 20+ 0.75Y

Y= 20+ 0.75Y+ 20

Y= 40+ 0.75Y

0.25Y = 40

Y= 160

b) When G= 20

Y= C+I+ G

Y= 20+ 0.75Y+ 20+ 20

Y=60+ 0.75Y

0.25Y= 60

Y= 240

d) when tax = 25

Y= C+I+G

C= 20+ 0.75(Y-25)

Y=20+ 0.75(Y-25)+ 20+20

Y= 60+ 0.75Y-18.75

0.25Y= 41.25

Y= 165